# Deep Reinforcement Learning-Based Active Flow Control of an Elliptical Cylinder: Transitioning from an Elliptical Cylinder to a Circular Cylinder and a Flat Plate

### v1

@@ II.1 Numerical simulation
Taking the case with an $Ar$ of 1 as an example, this study explores mesh convergence and validates the numerical method utilized. Tab. 1 presents the simulation results using three meshes with differing resolutions, enabling a comparative analysis of the maximum drag coefficient $C_{D,\max}$, the mean of the maximum drag coefficient $C_{D,\text{mean}}$, the maximum lift coefficient $C_{L,\max}$, and the $St$. Furthermore, the results of the numerical simulations are compared with the findings of Rabault et al. Rabault et al. (2019) and Schäfer et al. Schäfer et al. (1996) for validation. The simulations conducted using a coarse mesh revealed significant deviations in the drag and lift coefficients when compared to the results by Rabault et al. Rabault et al. (2019), questioning the accuracy of such numerical simulations. On the other hand, the fine mesh, characterized by a high density of grid points, considerably slows down the training speed for reinforcement learning, presenting a challenge for efficient simulation execution. Our aim is to perform numerical simulations with the least number of mesh elements without compromising on computational accuracy. Consequently, the discretization strategy of the main mesh was adopted for subsequent numerical simulations and reinforcement learning training. This strategy not only satisfies the computational accuracy requirements but also keeps the mesh element count to a manageable level. Comparison of the computational results from the main mesh with those obtained by Rabault et al. Rabault et al. (2019). shows very little difference, indicating a high level of accuracy. Furthermore, the accuracy of the main mesh’s results lies within the upper and lower bounds provided by standard benchmarksSchäfer et al. (1996), confirming that the computational precision of the main mesh is adequate for further training purposes. The above analysis ensures the best balance between numerical accuracy and computational efficiency. The methodology for examining mesh dependency for ellipses with $Ar$ other than 1 is consistent with the procedure established for $Ar=1$.
@@
Coarse 10540 3.242 3.224 1.052 0.304
$Ar=1$ Medium 18484 3.225 3.205 0.990 0.300
Fine 25624 3.228 3.207 0.992 0.301
@@ II.2 Deep Reinforcement learning
Deep learning, a subset of machine learning, employs multi-layer artificial neural networks to model and learn intricate nonlinear relationships. Fig. 3(a) demonstrates how deep learning captures the representation of input data across multiple layers of a neural network, including the input layer, several hidden layers, and the output layer. Each layer is composed of numerous neural units, with information transmitted between nodes via connection weights. The fundamental concept behind deep learning is its ability to autonomously learn data representations through successive layers. Utilizing the backpropagation algorithm, the network autonomously discerns the transformation process from raw data to sophisticated, high-level abstract features, eliminating the need for manual feature extraction.
(a) DL
(b) RL
(c) DRL
Figure 3: The basic architecture of deep learning, reinforcement learning, and deep reinforcement learning.
@@
In the domain of Reinforcement Learning, the interaction between an agent and its environment is the core mechanism by which learning is facilitated. As illustrated in Figure 3(b), this interaction proceeds in a cyclical pattern. At each timestep $t$, the agent observes the current state $s_{t}$ of the environment, which contains all the relevant information necessary for making decisions. Based on this observation, the agent selects and executes an action $a_{t}$, according to its policy $\pi$. This policy, which is essentially a strategy that maps states to actions, may function deterministically or probabilistically. After the action is executed, the environment transitions to a new state $s_{t+1}$ and awards the agent with an immediate reward $r_{t+1}$. This reward, quantified as a numerical value, evaluates the action’s efficacy in progressing towards the agent’s objectives. Subsequently, the agent assimilates this feedback—comprised of both the reward and the new state—to refine its policy, employing RL algorithms designed to optimize the accumulation of long-term rewards.
@@
Deep learning, a specialized subset of machine learning, excels at identifying patterns in extensive datasets through its multi-layer neural network architecture. This approach enables computers to derive complex representations directly from raw data, leading to revolutionary advances in fields like image recognition, natural language processing, and predictive analytics. DRL marries the potent perceptual capabilities of deep learning with the decision-making and optimization capabilities of reinforcement learning, effectively overcoming their individual limitations. Deep learning addresses the challenge of processing original, high-dimensional input data, which is problematic for traditional reinforcement learning, while reinforcement learning provides a structured decision-making framework and a goal-oriented learning mechanism that enhances the capabilities of deep learning. As depicted in Figure 3(c), a deep learning network transforms the intricate state of the environment into a compressed and efficient feature representation. This transformation yields a more manageable and interpretable state space for reinforcement learning algorithms. In DRL, reinforcement learning leverages these high-level features to learn and optimize decision-making policies. These policies are typically parameterized by deep neural networks, implying that the selection of actions is influenced not only by the current observation of the environmental state but also by the learned weights within the network.
II.3 PPO algorithm
PPO is an advanced DRL algorithm suitable for solving reinforcement learning problems in discrete and continuous action spaces. The core idea is to improve performance by maximizing approximate policy optimization while maintaining the stability of policy updates. In PPO, the agent interacts with the environment, collects empirical data, and uses this data to construct an objective function. The objective function takes into account dominance estimates from empirical data and is used to measure the direction of policy improvement. In the PPO algorithm, the agent interacts with the environment to generate a series of trajectories ($\tau$). A trajectory is a sequence of state-action pairs produced by the agent following the current policy interacting with the environment until reaching a terminal state or a predetermined maximum number of steps. Symbolically, a trajectory $\tau$ can be represented as:
@@ II.3 PPO algorithm
In summary, the PPO algorithm presents a robust framework for the training of deep reinforcement learning agents. It operates through an iterative process of updating policy parameters aimed at maximizing the expected cumulative reward. Crucially, PPO maintains stability and mitigates the risk of substantial policy deviations. Employing policy gradients along with clipped surrogate objectives, PPO ensures effective and stable training across a diverse array of reinforcement learning tasks. This methodology not only enhances the efficiency of learning but also contributes significantly to the consistency and reliability of the training process.
@@ II.4 DRL-Enhanced Active Flow Control
This section outlines the design of a framework utilizing DRL algorithms to address active flow control problems. Firstly, it is necessary to define two fundamental elements in DRL: the agent and the environment. The agent can be designed as algorithms suitable for solving continuous action control problems, such as PPO agent or SAC agent. The CFD simulation of the flow around an elliptical cylinder is chosen as the environment for the agent to interact with. Rabault et al.Rabault et al. presented the main structure of the custom environment class, encouraging readers to delve into the detailed content. This framework design applies deep reinforcement learning methods to solve the active flow control problem, enabling the agent to continuously learn and optimize its behavior through interaction with the environment to achieve the control goal. Fig. 4 illustrates the fundamental framework of DRL, the computational process of a single episode during training, and the formulas utilized.
@@
Figure 4: DRL leverages the concurrent execution of multiple environments within a parallel computational framework. Each iteration consists of two distinct phases: exploration and optimization. During the exploration phase, the agent engages with the environment via its current policy $\pi$, executing actions $a_{t}$, receiving states $s_{t}$ and rewards $r_{t}$, and noting terminal states. Upon the conclusion of the iteration, the parallel computation across $n$ environments yields a set of trajectories $\{\tau_{i}\}_{i=1}^{n}$, with each trajectory $\tau_{i}$ comprising the sequence of states, actions, and rewards as $\tau_{i}=\{(s_{0},a_{0}),(s_{1},a_{1},r_{1}),\ldots,(s_{T},a_{T},r_{T})\}$, where $(s_{t},a_{t},r_{t})$ represents the state, action , and reward at time-step $t$, respectively, for the $i$-th environment. In the optimization phase, the agent samples a minibatch of trajectories from the experience replay buffer to compute the policy gradient $\nabla_{\theta}J(\pi_{\theta})$. The parameters $\theta$ of the policy network are updated using gradient ascent, i.e., $\theta\leftarrow\theta+\alpha\nabla_{\theta}J(\pi_{\theta})$, where $\alpha$ denotes the learning rate. This process iterates, with the agent deploying the updated policy $\pi_{\theta^{\prime}}$ to interact with the environment, computing the advantage function $A(s_{t},a_{t})$, and recalculating the policy gradient for subsequent updates of $\theta$. Through iterative refinement, the policy network incrementally enhances its action selection strategy within the environment. The formula delineates the update rule applied to the policy network during the optimization phase.
@@
  • State $s_{t}$: Traditional state spaces are often too expansive for effective policy learning, requiring operations in partially observable environments. As a result, an agent’s performance heavily relies on the quality and relevance of the data it can observe. The impact of partial observability of states $s_{t}$ on control results was mentioned in Viquerat et al.Viquerat et al. (2022)’s research.Integrating instantaneous velocity or pressure information from CFD simulations to characterize the environmental state $s_{t}$, the DRL agent can effectively perceive and analyze the current flow conditions. To achieve the control objectives of reducing drag and lift around the elliptical cylinder and controlling the shedding of vortices in the wake field, multiple probes are strategically placed around the elliptical cylinder and within its wake field. This method is advantageous for gathering comprehensive data pertaining to the flow dynamics around the cylinder and subsequent wake characteristics. The design of probes placement in this study was based on the design principles articulated by Rabault et al.Rabault et al. (2019), which advocate strategic positioning of probes around the cylinder and within the wake region of the cylinder.In their studies, Jia and XuJia and Xu (2024a) and Wang et al.Wang et al. (2022b) all referred to the probes arrangement scheme proposed by Rabault et al.Rabault et al. (2019).
@@
  • Action $a_{t}$ : The control action involves the blowing and suction of two synthetic jets on the elliptical cylinder. The flow rates of the two synthetic jets are constrained such that $Q_{\text{lower}}+Q_{\text{upper}}=0$. Therefore, the RL agent manipulates the velocity of the upper synthetic jets, while the lower synthetic jets has the same velocity but in the opposite direction.The velocity of the synthetic jets is set to be less than 2% of the inlet mean velocity. This restriction is imposed to prevent the agent from taking shortcuts and using excessively high jet velocities to achieve drag reduction, which contradicts the fundamental principle of active flow control: energy efficiency. Furthermore, each control action lasts for 50 time steps, termed as a viscous action. Allowing faster actions could harm overall performance as they need time to influence the environment. Ensuring smooth transitions between current and updated actions is crucial. Let $a_{t}$ and $a_{t+1}$ represent the action magnitudes at consecutive time steps. To ensure a gradual transition between these actions and mitigate abrupt changes in jet velocity, a smoothing function $S$ can be defined mathematically. The function $S$ interpolates between the previous action magnitude $V_{\Gamma_{1},T_{i-1}}$ and the target action magnitude $a$ at time step $i$, resulting in an updated value $V_{\Gamma_{1},T_{i}}$. This smoothing function can be expressed as:
@@ II.5 High-Performance Computing in DRL and CFD
The DRL training and numerical simulation in this study were conducted on a high-performance computing(HPC) system equipped with an Intel® Xeon® Platinum 8358 CPU, operating at 2.60GHz and comprising a total of 64 cores distributed evenly across two sockets, each containing 32 cores. Numerical simulations were performed using the eighth version of the open-source software platform, OpenFOAM® version 8.Jasak et al. (2007) The deployment of deep reinforcement learning algorithms was based on the Tensorforce library.Schaarschmidt, Kuhnle, and Fricke (2017) The ANN definitions and gradient descent algorithms were facilitated using the TensorFlow open-source library.Abadi et al. (2016) For DRL implementation, agent was loaded using the Tensorforce platform with the PPO algorithm, and the environment was constructed using the Gym interface.Brockman et al. (2016) Fig. 4 illustrates the use of CFD calculations as the environment for DRL training and shows the effects of using 60 parallel CFD environments to accelerate the training process. This parallel strategy was inspired by the research of Rabault and KuhnleRabault and Kuhnle (2019a) and Wang et al.Wang et al. (2022b), and also incorporated insights from the work of Wang et al.Wang et al. (2022b) For a more comprehensive understanding, readers are encouraged to consult their works and the references therein.
@@ III.1 Baseline Flow
Figure 5: Instantaneous Vorticity Contours for Flow Around an Elliptic Cylinder at $Re=100$, Varying from $Ar=2$ to $Ar=0$. Purple and orange indicate positive and negative vorticity, respectively. In all the depicted vorticity contours, purple represents the clockwise vortices shed from the upper side of the elliptical cylinder, while orange represents the counterclockwise vortices shed from the lower side.
Fig. 5 presents the vortex shedding patterns in the wake region formed around elliptical cylinders with varying $Ar$. The subfigures (a) to (j) depict the vortex structures that have reached a fully developed state, illustrating the progression from an aspect ratio of 0 to 2. At a Reynolds number of 100, around elliptical cylinders with $Ar$ varying from 0 to 2, persistent unstable vortices are observed at the computational domain’s outlet, indicating the continuation of fluid instability along the outflow direction. However, as the fluid progresses downstream, there is a diminishing trend in the peak vortex intensity. This attenuation of vorticity is predominantly attributed to viscous diffusion effects prevalent in low Reynolds number regimes. This is consistent with the study by Johnson, Thompson, and HouriganJohnson, Thompson, and Hourigan (2004). Between $Ar$ of 2 and 1.25, in addition to the recirculation bubbles downstream of the cylinder, two additional recirculation bubbles form upstream near the wall. As the $Ar$ decreases further, the confined flow exhibits only one recirculation bubble. Similar to observations by Sahin and OwensSahin and Owens (2004), under the influence of blockage ratio and Reynolds number, two additional recirculation bubbles appear on the solid sidewalls of the computational domain. At specific Reynolds numbers, this will lead to another Hopf bifurcation at $Re_{\text{c2}}$.
@@
For an elliptical cylinder with an $Ar=2$, the wake field does not exhibit vortex shedding, and the zero vorticity contour lines show slight fluctuations. This suggests that despite the overall stability of the wake, there is still minor flow instability. As the $Ar$ decreases to 1.75 and 1.5, the wake field begins to exhibit more pronounced vortex shedding phenomena. The vortices become more compact, and there is increased interaction between positive and negative vortices through shearing, stretching, and detachment processes. As the $Ar$ decreases to 1, transforming the cylinder into a perfect circle, the vortex structure becomes more regular, forming a typical von Kármán vortex street. This "standard aligned BvK street" is characterized by a series of regularly spaced vortices that alternately form on both sides of the cylinder. As the $Ar$ decreases from $Ar=0.75$ to $Ar=0.25$, the recirculation area gradually diminishes, leading to a more compact vortex structure and an increase in shedding frequency. At $Ar=0.1$, for extremely low $Ar$, the recirculation area further decreases, resulting in an increase in both the frequency and intensity of vortex shedding. At $Ar=0$, when the cylinder degenerates into a flat plate, particularly in the zero-vortex region behind the ellipse, the acceleration of vortex shedding becomes more pronounced. This indicates an increase in flow instability and irregularity in vortex size and shape.
Fluctuation velocity
Strykowski and SreenivasanStrykowski and Sreenivasan (1990) demonstrated that the introduction of minor disturbances in a targeted region can alter vortex formation, effectively reducing or controlling vortex shedding. Similarly, our study employs small perturbations to disturb the flow field with the aim of decreasing drag and suppressing vortex shedding. Furthermore, the judicious incorporation of fluid physics knowledge within the control agents can streamline their architecture without compromising performance, as exemplified by Pastoor et al.Pastoor et al. (2008), where phase control in drag reduction of a bluff body is primarily founded upon the understanding of decoupled shear layer development and wake processes. In this study, the strategic placement of probes was thoughtfully engineered to align with the dynamic and controllable aspects of the flow field. Particular attention was paid to its inherent instability and pulsatility, ensuring that the probes could effectively capture the critical fluctuations necessary for a comprehensive analysis of the flow dynamics. The research conducted by Protas and WesfreidProtas and Wesfreid (2002) explores the dynamics of the total drag coefficient, $C_{D}$, which is defined by the equation $C_{D}=C_{D}^{b}+C_{D}^{0}$. Here, $C_{D}^{b}$ represents the base drag coefficient, a static value consistent across specific $Re$, encapsulating the inherent drag associated with the baseline flow. Conversely, the correctable component $C_{D}^{0}$ reflects the influence of oscillatory flow phenomena on the drag coefficient, providing an avenue for adjustment through strategic manipulation of Bénard–von Kármán vortices. In our study, the instantaneous velocity field is segmented into the mean velocity field and the fluctuating velocity field. We hypothesize that probes equipped to detect these pulsatile and oscillatory dynamics can leverage targeted minor disturbances to effectively neutralize these instabilities. Such interventions are anticipated to facilitate the suppression of vortex shedding, thereby enhancing the control of flow behaviors.
@@
Figure 6: For the case of $Ar=1$, the distribution of probes is depicted across various velocity fields: (a) Instantaneous velocity, (b) Fluctuating velocity.
We delineates the instantaneous and fluctuating velocity fields behind an elliptical cylinder during full vortex development, as illustrated in Fig. 6. The strategic placement of probes is predicated on the objectives of reducing drag and lift around the elliptical cylinder while simultaneously suppressing vortex shedding. Probes are positioned around and downstream of the elliptical cylinder, specifically within the vortex shedding region. The probes around the elliptical cylinder aim to diminish the drag and lift coefficients integral to our control objectives. Downstream, probes are strategically deployed in the wake to facilitate the extraction of flow field information from the regions of maximum fluctuating velocity, where intense turbulent motion typically occurs due to vortex instability and enhanced turbulent energy. The probe placement in the cylinder’s wake encompasses the entire recirculation area, which is critical for capturing the most unstable and intensely turbulent information in the wake region. This configuration plays a significant role in reducing the fluctuating components of the drag coefficient, suppressing vortex shedding, and stabilizing the wake flow. In Fig. 6, an ellipse with an $Ar$ of 1 serves as the prototype. The methodology for designing probe placements for other $Ar$ adheres to the same principles established for $Ar=1$.

### v2

@@ II.1 Numerical simulation
Taking the case with an $Ar$ of 1 as an example, this study explores mesh convergence and validates the numerical method utilized. Tab. 1 presents the simulation results using three meshes with differing resolutions, enabling a comparative analysis of the maximum drag coefficient $C_{D,\max}$, the mean of the maximum drag coefficient $C_{D,\text{mean}}$, the maximum lift coefficient $C_{L,\max}$, and the $St$. Furthermore, the results of the numerical simulations are compared with the findings of Rabault et al. Rabault et al. (2019) and Schäfer et al. Schäfer et al. (1996) for validation. The simulations conducted using a coarse mesh revealed significant deviations in the drag and lift coefficients when compared to the results by Rabault et al. Rabault et al. (2019), questioning the accuracy of such numerical simulations. On the other hand, the fine mesh, characterized by a high density of grid points, considerably slows down the training speed for reinforcement learning, presenting a challenge for efficient simulation execution. Our aim is to perform numerical simulations with the least number of mesh cells without compromising on computational accuracy. Consequently, the discretization strategy of the main mesh was adopted for subsequent numerical simulations and reinforcement learning training. This strategy not only satisfies the computational accuracy requirements but also keeps the mesh element count to a manageable level. Comparison of the computational results from the main mesh with those obtained by Rabault et al. Rabault et al. (2019). shows very little difference, indicating a high level of accuracy. Furthermore, the accuracy of the main mesh’s results lies within the upper and lower bounds provided by standard benchmarksSchäfer et al. (1996), confirming that the computational precision of the main mesh is adequate for further training purposes. The above analysis ensures the best balance between numerical accuracy and computational efficiency. The methodology for examining mesh dependency for ellipses with $Ar$ other than 1 is consistent with the procedure established for $Ar=1$.
@@
Coarse 10,540 3.242 3.224 1.052 0.304
$Ar=1$ Medium 18,484 3.225 3.205 0.990 0.300
Fine 25,624 3.228 3.207 0.992 0.301
@@ II.2 Deep Reinforcement learning
Deep learning, a subset of machine learning, employs multi-layer artificial neural networks to model and learn intricate nonlinear relationships. Fig. 3 demonstrates how deep learning captures the representation of input data across multiple layers of a neural network, including the input layer, several hidden layers, and the output layer. Each layer is composed of numerous neural units, with information transmitted between nodes via connection weights. The fundamental concept behind deep learning is its ability to autonomously learn data representations through successive layers. Utilizing the backpropagation algorithm, the network autonomously discerns the transformation process from raw data to sophisticated, high-level abstract features, eliminating the need for manual feature extraction.
Figure 3: DRL combines deep learning’s ability to interpret complex data with reinforcement learning’s strategy to make informed decisions.
@@
In the domain of Reinforcement Learning, the interaction between an agent and its environment is the core mechanism by which learning is facilitated. As illustrated in Figure 3, this interaction proceeds in a cyclical pattern. At each timestep $t$, the agent observes the current state $s_{t}$ of the environment, which contains all the relevant information necessary for making decisions. Based on this observation, the agent selects and executes an action $a_{t}$, according to its policy $\pi$. This policy, which is essentially a strategy that maps states to actions, may function deterministically or probabilistically. After the action is executed, the environment transitions to a new state $s_{t+1}$ and awards the agent with an immediate reward $r_{t+1}$. This reward, quantified as a numerical value, evaluates the action’s efficacy in progressing towards the agent’s objectives. Subsequently, the agent assimilates this feedback—comprised of both the reward and the new state—to refine its policy, employing RL algorithms designed to optimize the accumulation of long-term rewards.
@@
Deep learning, a specialized subset of machine learning, excels at identifying patterns in extensive datasets through its multi-layer neural network architecture. This approach enables computers to derive complex representations directly from raw data, leading to revolutionary advances in fields like image recognition, natural language processing, and predictive analytics. DRL marries the potent perceptual capabilities of deep learning with the decision-making and optimization capabilities of reinforcement learning, effectively overcoming their individual limitations. Deep learning addresses the challenge of processing original, high-dimensional input data, which is problematic for traditional reinforcement learning, while reinforcement learning provides a structured decision-making framework and a goal-oriented learning mechanism that enhances the capabilities of deep learning. As depicted in Figure 3, a deep learning network transforms the intricate state of the environment into a compressed and efficient feature representation. This transformation yields a more manageable and interpretable state space for reinforcement learning algorithms. In DRL, reinforcement learning leverages these high-level features to learn and optimize decision-making policies. These policies are typically parameterized by deep neural networks, implying that the selection of actions is influenced not only by the current observation of the environmental state but also by the learned weights within the network.
II.3 Proximal Policy Optimization algorithm
PPO is an advanced DRL algorithm suitable for addressing reinforcement learning challenges within both discrete and continuous action spaces. The fundamental concept of PPO revolves around enhancing performance by maximizing a surrogate objective function while ensuring the stability of policy updates. In PPO, an agent interacts with an environment to gather experience data, which is then utilized to construct the objective function. This function incorporates dominant estimates from the experience data to gauge the direction of policy improvement. During these interactions, the agent generates a sequence of trajectories ($\tau$). A trajectory consists of a series of state-action pairs produced as the agent operates under the current policy, continuing until a terminal state is reached or a predetermined maximum number of steps is exceeded. Symbolically, a trajectory $\tau$ can be represented as follows:
@@ II.3 Proximal Policy Optimization algorithm
In summary, the PPO algorithm presents a robust framework for the training of DRL agents. It operates through an iterative process of updating policy parameters aimed at maximizing the expected cumulative reward. Crucially, PPO maintains stability and mitigates the risk of substantial policy deviations. Employing policy gradients along with clipped surrogate objectives, PPO ensures effective and stable training across a diverse array of reinforcement learning tasks. This methodology not only enhances the efficiency of learning but also contributes significantly to the consistency and reliability of the training process.
@@ II.4 DRL-Enhanced Active Flow Control
This section outlines the design of a framework utilizing DRL algorithms to address AFC problems. Firstly, it is necessary to define two fundamental elements in DRL: the agent and the environment. The agent can be designed as algorithms suitable for solving continuous action control problems, such as PPO agent or SAC agent. The CFD simulation of the flow around an elliptical cylinder is chosen as the environment for the agent to interact with. Rabault et al.Rabault et al. (2020) presented the main structure of the custom environment class, encouraging readers to delve into the detailed content. This framework design applies DRL methods to solve the AFC problem, enabling the agent to continuously learn and optimize its behavior through interaction with the environment to achieve the control goal. Fig. 4 illustrates the fundamental framework of DRL, the computational process of a single episode during training, and the formulas utilized.
@@
Figure 4: DRL leverages the concurrent execution of multiple environments within a parallel computational framework. Each iteration consists of two distinct phases: exploration and optimization. During the exploration phase, the agent engages with the environment via its current policy $\pi$, executing actions $a_{t}$, receiving states $s_{t}$ and rewards $r_{t}$, and noting terminal states. Upon the result of the iteration, the parallel computation across $n$ environments yields a set of trajectories $\{\tau_{i}\}_{i=1}^{n}$, with each trajectory $\tau_{i}$ comprising the sequence of states, actions, and rewards as $\tau_{i}=\{(s_{0},a_{0}),(s_{1},a_{1},r_{1}),\ldots,(s_{T},a_{T},r_{T})\}$, where $(s_{t},a_{t},r_{t})$ represents the state, action , and reward at time-step $t$, respectively, for the $i$-th environment. In the optimization phase, the agent samples a minibatch of trajectories from the experience replay buffer to compute the policy gradient $\nabla_{\theta}J(\pi_{\theta})$. The parameters $\theta$ of the policy network are updated using gradient ascent, i.e., $\theta\leftarrow\theta+\alpha\nabla_{\theta}J(\pi_{\theta})$, where $\alpha$ denotes the learning rate. This process iterates, with the agent deploying the updated policy $\pi_{\theta^{\prime}}$ to interact with the environment, computing the advantage function $A(s_{t},a_{t})$, and recalculating the policy gradient for subsequent updates of $\theta$. Through iterative refinement, the policy network incrementally enhances its action selection strategy within the environment.
@@
  • State $s_{t}$: Traditional state spaces are often too expansive for effective policy learning, requiring operations in partially observable environments. As a result, an agent’s performance heavily relies on the quality and relevance of the data it can observe. The impact of partial observability of states $s_{t}$ on control results was mentioned in Viquerat et al.Viquerat et al. (2022)’s research.Integrating instantaneous velocity or pressure information from CFD simulations to characterize the environmental state $s_{t}$, the DRL agent can effectively perceive and analyze the current flow conditions. To achieve the control objectives of reducing drag and lift around the elliptical cylinder and controlling the shedding of vortices in the wake field, multiple probes are strategically placed around the elliptical cylinder and within its wake field. This method is advantageous for gathering comprehensive data pertaining to the flow dynamics around the cylinder and subsequent wake characteristics. The design of probes placement in this study was based on the design principles articulated by Rabault et al.Rabault et al. (2019), which advocate strategic positioning of probes around the cylinder and within the wake region of the cylinder.In their studies, Wang and XuWang and Xu (2024) and Wang et al.Wang et al. (2022a) all referred to the probes arrangement scheme proposed by Rabault et al.Rabault et al. (2019).
@@
  • Action $a_{t}$ : The control action involves the blowing and suction of two synthetic jets on the elliptical cylinder. The flow rates of the two synthetic jets are constrained such that $Q_{\text{lower}}+Q_{\text{upper}}=0$. Therefore, the RL agent manipulates the velocity of the upper synthetic jets, while the lower synthetic jets has the same velocity but in the opposite direction. The velocity of the synthetic jets is set to be less than 2% of the inlet mean velocity. This restriction is imposed to prevent the agent from taking shortcuts and using excessively high jet velocities to achieve drag reduction, which contradicts the fundamental principle of AFC: energy efficiency. Furthermore, each control action lasts for 50 time steps, termed as a viscous action. Allowing faster actions could harm overall performance as they need time to influence the environment. Ensuring smooth transitions between current and updated actions is crucial. Let $a_{t}$ and $a_{t+1}$ represent the action magnitudes at consecutive time steps. To ensure a gradual transition between these actions and mitigate abrupt changes in jet velocity, a smoothing function $S$ can be defined mathematically. The function $S$ interpolates between the previous action magnitude $V_{\Gamma_{1},T_{i-1}}$ and the target action magnitude $a$ at time step $i$, resulting in an updated value $V_{\Gamma_{1},T_{i}}$. This smoothing function can be expressed as:
@@ II.5 High-Performance Computing in DRL and CFD
The DRL training and numerical simulation in this study were conducted on a high-performance computing (HPC) system equipped with an Intel® Xeon® Platinum 8358 CPU, operating at 2.60 GHz and comprising a total of 64 cores distributed evenly across two sockets, each containing 32 cores. Numerical simulations were performed using the eighth version of the open-source software platform, OpenFOAM® version 8.Jasak et al. (2007) The deployment of DRL algorithms was based on the Tensorforce library.Schaarschmidt et al. (2018) The ANN definitions and gradient descent algorithms were facilitated using the TensorFlow open-source library.Abadi et al. (2016) For DRL implementation, agent was loaded using the Tensorforce platform with the PPO algorithm, and the environment was constructed using the Gym interface.Brockman et al. (2016) Fig. 4 illustrates the use of CFD calculations as the environment for DRL training and shows the effects of using 60 parallel CFD environments to accelerate the training process. This parallel strategy was inspired by the research of Rabault and KuhnleRabault and Kuhnle (2019a) and Wang et al.Wang et al. (2022a), and also incorporated insights from the work of Wang et al.Wang et al. (2022a) For a more comprehensive understanding, readers are encouraged to consult their works and the references therein.
@@ III.1 Baseline Flow
Fig. 5 depicts the vortex shedding patterns within the wake field of elliptical cylinders with $Ar$ from 0 to 2. Starting from Fig. 5(a) depicting an elliptical cylinder with $Ar=2$, the wake field does not exhibit vortex shedding. This is expected as the elongated elliptical cylinder results in weak adverse pressure gradients that allow the boundary layers to stay attached for a prolonged distance. The zero vorticity contour lines in the wake extend all the way to the outlet with slight fluctuations. This suggests that the wake is largely stable, with minor flow instability due to shear in the flow. It is worth highlighting that this test case of $Ar=2$ serves as a special verification study for subsequent DRL control experiments. For a stable flow field with no vortex shedding, we expect the DRL-based flow control algorithm to conform to the underlying flow physics and to not introduce any actuation since the current baseline is essentially the optimal flow configuration. To the best of our knowledge, this sanity test that verifies minimal DRL-based control is often overlooked in literature. We include this to showcase the robustness of DRL-based control across a spectrum of flow physics. We delay the discussion of DRL control to Sec. III.2.
@@
Figure 5: Contours of instantaneous vorticity for flow around an elliptic cylinder at $Re=100$, varying from $Ar=2$ to $Ar=0$. Purple and orange indicate positive and negative vorticity, respectively. In all the depicted vorticity contours, purple represents the clockwise vortices shed from the upper side of the elliptical cylinder, while orange represents the counterclockwise vortices shed from the lower side.
@@
As the $Ar$ decreases to 1.75, the recirculation region behind the cylinder starts to become unstable and begins to exhibit vortex shedding phenomena. As $Ar$ decreases further to 1.5 and 1.25, the vortices become more compact with increasing frequency, and there is increasing interaction between pairs of vortices through shearing, stretching, and detachment processes. It is also evident that the onset of vortex shedding shifts upstream with decreasing $Ar$. As the $Ar$ decreases to 1, which transforms the elliptical cylinder into a perfect circle, we observe the typical von Kármán vortex street reported in many literature. This "standard aligned BvK street" is characterized by a series of regularly spaced vortices that alternately form on both sides of the cylinder. As the $Ar$ decreases from $Ar=0.75$ to $Ar=0.25$, the recirculation area gradually diminishes, leading to a more compact vortex structure and an increase in shedding frequency. At $Ar=0.1$, for extremely low $Ar$, the recirculation area further decreases, resulting in an increase in both the frequency and intensity of vortex shedding. At $Ar=0$, when the cylinder degenerates into a vertical flat plate, flow separates right at the two sides of the flat plate and vortex shedding becomes more pronounced. This indicates an increase in flow instability and irregularity in vortex size and shape.
Across all test cases, persistent instabilities in the form of vortices are observed at the exit of the computational domain, indicating that the fluid flow remains unstable in the downstream direction. As the fluid advances downstream, there is a tendency for the peak vortex intensity to decrease. This attenuation of vorticity is primarily attributed to the viscous diffusion effects commonly present in low Reynolds number regions. This finding is consistent with the studies by Johnson, Thompson, and Hourigan Johnson, Thompson, and Hourigan (2004). Between $Ar$ of 2 and 1.25, in addition to the recirculation bubble downstream of the cylinder, two additional recirculation bubbles form near the upstream wall. As $Ar$ is further reduced, the confined flow exhibits only one recirculation bubble. This observation is similar to that of Sahin and Owens Sahin and Owens (2004), where two additional recirculation bubbles appear on the solid sidewalls of the computational domain under the influence of blockage ratio and Reynolds number. These phenomena in the wake field of elliptical cylinders represent some intriguing physical behaviors. This serves as a foundation for subsequent flow control studies.
Velocity fluctuation and design of probe layout
In addition to the baseline flow field, we analyze the fluctuating component of velocity, which serves as a good indicator to inform the location of probes during control. The judicious incorporation of fluid physics knowledge within the control agents can streamline their architecture without compromising performance, as exemplified by Pastoor et al.Pastoor et al. (2008), where phase control in drag reduction of a bluff body is primarily founded upon the understanding of decoupled shear layer development and wake processes. We also utilize the physical information from the wake field to guide the flow control strategies of the agents. This aspect is particularly reflected in the practical design of the probe positions. The study by Protas and Wesfreid Protas and Wesfreid (2002) elaborates on the adjustable component of the drag coefficient, $C_{D}^{0}$, which reflects the influence of oscillatory flow phenomena on the drag coefficient. This provides a pathway for strategic manipulation of the von Kármán vortex street to adjust the drag through targeted interventions. Building on this foundation, we tailor the physical information observed by the agents to specifically face the factors contributing to instability in the wake field.
@@
Figure 6: The distribution of probes within the fluctuating velocity magnitude fields for different $Ar$ : (a) $Ar=0.25$; (b) $Ar=0.1$; (c) $Ar=0$.
As depicted in Fig. 6, the fluctuating magnitude velocity fields around elliptical cylinders with $Ar$ of 0, 0.1, and 0.25 are illustrated. The positions of the black dots on the diagram represent the locations of the probes. These probes are strategically placed around the elliptical cylinder and in the downstream wake region. On one hand, the arrangement of probes around the elliptical cylinder is designed to capture physical information pertinent to the lift and drag forces, which are critical components of our control objectives. On the other hand, probes within the wake of the elliptical cylinder are positioned to encompass areas of maximum velocity fluctuations, facilitating the extraction of flow field data from regions where velocity fluctuations and vortex instabilities are most intense. The extraction of physical information from these locations enables the agents to acutely observe changes in the flow field. In Fig. 6, three tests with different $Ar$ are shown as examples. The configurations for other $Ar$ are fundamentally similar and are not elaborated further.